# Privately Estimating Black-Box Statistics

### v1

@@ 5 Lower Bound
Our lower bound in Theorem 5.1 assumes the algorithm $M$ has an infinite range $\mathcal{Y}=\mathbb{Z}$. In contrast, our algorithm (Theorem 4.3) assumes a finite range $\mathcal{Y}$ and has a dependency on its size, namely $t=\frac{1}{\varepsilon}\log(1/\delta)\exp(O(\log^{*}|\mathcal{Y}|))$. However, the proof of Theorem 5.1 assumes the algorithm $M$ has range $\mathcal{Y}=[\ell^{2}]$ (and we take $\ell\to\infty$). Thus our lower bound could be extended to the setting with a finite range.
@@ 6.1 Interpretation
By varying the parameter $m$ (with $n\geq m\geq t\approx\Theta(\frac{1}{\varepsilon}\log(1/\delta))$) our result interpolates between sample-and-aggregate [NRS07] and more recent results [FDY22, LRSS25]. While we believe that the entire tradeoff curve is interesting, we point out a few illustrative values along the curve in Table 6.1 to aid understanding.
@@
2 $\frac{2n}{t+2}$ $\leq\frac{(t+2)(t+1)}{2}=O(t^{2})$ (3.29,3.30)
3 $\frac{cn}{t+c}$ $\leq{t+c\choose c}=O(t^{c})$ $c\in\mathbb{N}$
@@
5 ntn-t ${n\choose t}=O(n/t)^{t}$ [[, Cf.]]FangDY22,DPorg-down-sensitivity,LRSS25
@@
Table 6.1: Example parameter choices for Theorem $1.$ This shows the tradeoff between the number of evaluations of the function $k$ (i.e., oracle complexity) and the size of the subsets on which we evaluate it (which determines statistical efficiency). Here $t=\frac{1}{\varepsilon}\log(1/\delta)\exp(O(\log^{*}|\mathcal{Y}|))$ depends on the privacy parameters $\varepsilon,\delta>0$ and the size of the range $\mathcal{Y}$ of the function.
@@
6.2 Limitations & Futher Work
@@
Choosing the subsets amounts to generating a covering design. As mentioned in Section 3.4, we do not have general-purpose optimal existential results for covering designs, let alone efficient constructions. (Although there are a lot of special-purpose constructions in the literature.) However, a (slightly suboptimal) covering design can be constructed (with high probability) by simply taking enough random subsets of the appropriate size (see the proof of Proposition 3.7 and Footnote 6). We must account for the failure probability, but otherwise this does give an efficient method for choosing the subsets. [^10] Verifying that a given collection of sets is a covering design is co-NP-hard. Thus this failure probability cannot easily be “checked away.” However, we remark that the privacy guarantee of our algorithm still holds if the sets do not form a covering design; namely Lemma 4.1 holds for any choice of subsets $S_{1},\cdots,S_{k}\subseteq[n]$.
@@
In the black-box setting the function $f$ is indeed arbitrary. The input $x$ is i.i.d., but from an arbitrary distribution. However, the sets $S_{1},\cdots,S_{n}$ are not arbitrary; our algorithm can choose them, subject only to the constraint that they form a covering design. This points to an avenue for making the algorithm computationally efficient – construct a covering design with additional structural properties that make the decision problem in Equation 6.3 easy. We formulate this as an open problem:
@@ A.4 Noisy Binary Search
We can find an appropriate index $i$ via binary search using $\log_{2}|\mathcal{Y}|$ steps, each time comparing the function value $\ell(x,y_{i})$ to the target value $\tau$. To ensure concentrated differential privacy, we add Gaussian noise $\mathcal{N}(0,\sigma^{2})$ to the function value $\ell(x,y_{i})$. This means that each comparison may be incorrect. The parameter $\eta>0$ allows us to tolerate some amount of error. By composition over $\log_{2}|\mathcal{Y}|$ steps, the scale of the noise grows as $\sigma=O(\sqrt{\log|\mathcal{Y}|})$ Naïvely, we must take a union bound over all $O(\log|\mathcal{Y}|)$ steps, which adds a $\sqrt{\log\log|\mathcal{Y}|}$ factor to our error guarantee – i.e., $\eta=O(\sigma\cdot\sqrt{\log\log|\mathcal{Y}|})=O(\sqrt{\log|\mathcal{Y}|\cdot\log\log|\mathcal{Y}|})$. (This is because the maximum of $k$ Gaussians is $\sqrt{\log k}$ standard deviations above the mean.) Fortunately, we can remove this asymptotic factor.
@@
We can simplify the bound to $q\leq O(\log(m/\beta)/\gamma^{2})$. [GP24] state a more general form of the result in which the target bias of $1/2$ can be set differently. The above special case of their result suffices for our application.

### v2

@@ 6.1 Interpretation
By varying the parameter $m$ (with $n\geq m\geq t\approx\Theta(\frac{1}{\varepsilon}\log(1/\delta))$) our result interpolates between sample-and-aggregate [NRS07] and more recent results [FDY22, LRSS25]. While we believe that the entire tradeoff curve is interesting, we point out a few illustrative values along the curve in Table 6.1 to aid understanding. We also plot some numerical values of the upper and lower bounds in Figure 2.
Figure 2 shows that the number of evaluations $k=C(n,m,t)$ can become extremely large. In particular, the practical parameter regime is where $m$ is close to $n$.
@@
2 $\frac{2n}{t+2}$ $\leq\frac{(t+2)(t+1)}{2}=O(t^{2})$ (3.14,3.15)
3 $\frac{cn}{t+c}$ $\leq O(\min\{e^{t+c}\cdot t\log m,t^{c}\})$ $c\in\mathbb{N}$
@@
5 ntn-t ${n\choose t}=O(n/t)^{t}$ [[, Cf.]]LRSS25
@@
Table 6.1: Example parameter choices for Theorem 1.1. This shows the tradeoff between the number of evaluations of the function $k$ (i.e., oracle complexity) and the size of the subsets on which we evaluate it (which determines statistical efficiency). Here $t=\frac{1}{\varepsilon}\log(1/\delta)\exp(O(\log^{*}|\mathcal{Y}|))$ depends on the privacy parameters $\varepsilon,\delta>0$ and the size of the range $\mathcal{Y}$ of the function.
@@
6.2 Limitations & Further Work
@@
Choosing the subsets amounts to generating a covering design. As mentioned in Section 3.4, we do not have general-purpose optimal existential results for covering designs, let alone efficient constructions. (Although there are a lot of special-purpose constructions in the literature.) However, a (slightly suboptimal) covering design can be constructed (with high probability) by simply taking enough random subsets of the appropriate size (see the proof of Proposition 3.7 and Footnote 17). We must account for the failure probability, but otherwise this does give an efficient method for choosing the subsets. [^11] Verifying that a given collection of sets is a covering design is co-NP-hard. Thus this failure probability cannot easily be “checked away.” However, we remark that the privacy guarantee of our algorithm still holds if the sets do not form a covering design; namely Lemma 4.1 holds for any choice of subsets $S_{1},\cdots,S_{k}\subseteq[n]$.
@@
In the black-box setting the function $f$ is indeed arbitrary. The input $x$ is i.i.d., but from an arbitrary distribution. Crucially, the sets $S_{1},\cdots,S_{n}$ are not arbitrary; our algorithm can choose them, subject only to the constraint that they form a covering design. This points to an avenue for making the algorithm computationally efficient – construct a covering design with additional structural properties that make the decision problem in Equation 6.3 easy. We formulate this precisely as an open problem:
@@
Finally, we remark that, although set cover is NP-complete, the runtime is exponential only in the size of the covering, which should be modest (at most $t$). That is, a naïve implementation of $\mathsf{Eval}$ would run in time $O(n^{v}\cdot k)$.
@@ A.4 Noisy Binary Search
We can find an appropriate index $i$ via binary search using $\log_{2}|\mathcal{Y}|$ steps, each time comparing the function value $\ell(x,y_{i})$ to the target value $\tau$. To ensure concentrated differential privacy, we add Gaussian noise $\mathcal{N}(0,\sigma^{2})$ to the function value $\ell(x,y_{i})$. This means that each comparison may be incorrect. The parameter $\eta>0$ allows us to tolerate some amount of error. By composition over $\log_{2}|\mathcal{Y}|$ steps, the scale of the noise grows as $\sigma=O(\sqrt{\log|\mathcal{Y}|})$. Naïvely, we must take a union bound over all $O(\log|\mathcal{Y}|)$ steps, which adds a $\sqrt{\log\log|\mathcal{Y}|}$ factor to our error guarantee – i.e., $\eta=O(\sigma\cdot\sqrt{\log\log|\mathcal{Y}|})=O(\sqrt{\log|\mathcal{Y}|\cdot\log\log|\mathcal{Y}|})$. (This is because the maximum of $k$ Gaussians is $\sqrt{\log k}$ standard deviations above the mean.) Fortunately, we can remove this asymptotic factor.
@@
We can simplify the bound (A.7) to $q\leq O(\log(m/\beta)/\gamma^{2})$. [GP24] state a more general form of the result in which the target bias of $1/2$ can be set differently. The above special case of their result suffices for our application.
@@
Appendix B Covering Designs
For completeness, we provide proofs of the bounds we use on the size of covering designs.
Proposition B.1 (Proposition 3.7).
For all $n,m,t\in\mathbb{N}$ with $n\geq m\geq t$ we have
$\frac{{n\choose t}}{{m\choose t}}\leq C(n,m,t)<\frac{{n\choose t}}{{m\choose t}}\left(1+\log{m\choose t}\right)+1.$ (B.1)
Proof.
The lower bound is due to [Sch64] and the upper bound is due to [ES74, Theorem 13.4]. Both proofs rely on the probabilistic method.
$C(n,m,t)\geq\frac{n}{m}C(n-1,m-1,t-1)\geq\frac{n}{m}\frac{n-1}{m-1}C(n-2,m-2,t-2)\geq\cdots\geq\prod_{i=0}^{t-1}\frac{n-i}{m-i}=\frac{{n\choose t}}{{m\choose t}}$. (B.2)
${\mathbb{P}\left[\not\exists i\in[k_{1}]~~T\subseteq S_{i}\right]}=\prod_{i\in[k_{1}]}\big(1-{\mathbb{P}\left[T\subseteq S_{i}\right]}\big)=\left(1-\frac{{n-t\choose m-t}}{{n\choose m}}\right)^{k_{1}}.$ (B.3)
${\mathbb{E}\left[K_{2}\right]}={n\choose t}{\mathbb{P}\left[\not\exists i\in[k_{1}]~~T\subseteq S_{i}\right]}={n\choose t}\left(1-\frac{{n-t\choose m-t}}{{n\choose m}}\right)^{k_{1}},$ (B.4)
$C(n,m,t)$ $\leq k_{1}+\left\lfloor{n\choose t}\left(1-\frac{{n-t\choose m-t}}{{n\choose m}}\right)^{k_{1}}\right\rfloor$ (B.5)
$C(n,m,t)\leq\left\lceil\frac{{n\choose t}}{{m\choose t}}\log{m\choose t}\right\rceil+\left\lfloor\frac{{n\choose t}}{{m\choose t}}\right\rfloor<\frac{{n\choose t}}{{m\choose t}}\left(\log{m\choose t}+1\right)+1,$ (B.6)
which is the desired result.181818[ES74] state the result without the +1+1 at the end, but it is not clear to us how they obtain this result. Their proof ignores the need to round $k_{1}$ to an integer and they leave fixing this issue as an exercise.
$C(n,m,t)\geq\left\lceil\frac{n}{m}C(n-1,m-1,t-1)\right\rceil$ for all $n\geq m\geq t\geq 1$. Equation B.2 then becomes
$C(n,m,t)\geq\left\lceil\frac{n}{m}C(n-1,m-1,t-1)\right\rceil\geq\left\lceil\frac{n}{m}\left\lceil\frac{n-1}{m-1}\left\lceil\frac{n-2}{m-2}\left\lceil\cdots\left\lceil\frac{n-t+1}{m-t+1}\right\rceil\cdots\right\rceil\right\rceil\right\rceil\right\rceil$ (B.7)
Corollary B.2 (Corollary 3.8).
For all $n,m,t\in\mathbb{N}$ with $n\geq m\geq t>1$ we have
$\left(\frac{n}{m}\right)^{t}\leq C(n,m,t)<\left(\frac{ne}{m}\right)^{t}\cdot\min\{1+m\log 2,1+t\log m\}+1.$ (B.8)
Proof.
Proposition 3.7 states that
$\frac{{n\choose t}}{{m\choose t}}\leq C(n,m,t)<\frac{{n\choose t}}{{m\choose t}}\left(1+\log{m\choose t}\right)+1.$ (B.9)
We simplify these bounds, starting with the upper bound. We use standard bounds on binomial coefficients: For all $m\geq t>1$, we have
$\left(\frac{m}{t}\right)^{t}\leq{m\choose t}\leq\min\left\{\left(\frac{em}{t}\right)^{t},m^{t},2^{m}\right\}.$ (B.10)
Thus $\log{m\choose t}\leq\min\{t+t\log(m/t),t\log m,m\log 2\}$ and
$\frac{{n\choose t}}{{m\choose t}}\leq\frac{\left(\frac{en}{t}\right)^{t}}{\left(\frac{m}{t}\right)^{t}}=\left(\frac{en}{m}\right)^{t},$ (B.11)
which gives the upper bound. For the lower bound, write
$\frac{{n\choose t}}{{m\choose t}}=\prod_{i=0}^{t-1}\frac{n-i}{m-i}=\exp\left(\sum_{i=0}^{t-1}f(i)\right),$ (B.12)
where we define $f(x)=\log\left(\frac{n-x}{m-x}\right)=\log(n-x)-\log(m-x)$. For $0\leq x<m\leq n$, we have $f^{\prime}(x)=\frac{-1}{n-x}-\frac{-1}{m-x}=\frac{n-m}{(m-x)(n-x)}\geq 0$. Thus $f$ is an increasing function on $[0,m)$. Using $\log(\frac{n}{m})=f(0)\leq f(i)\leq f(t-1)=\log(\frac{n-t+1}{m-t+1})$ gives
$\left(\frac{n}{m}\right)^{t}\leq\frac{{n\choose t}}{{m\choose t}}=\exp\left(\sum_{i=0}^{t-1}f(i)\right)\leq\left(\frac{n-t+1}{m-t+1}\right)^{t},$ (B.13)
which yields the lower bound. For $0\leq x<m\leq n$, we also have $f^{\prime\prime}(x)=\frac{-1}{(n-x)^{2}}-\frac{-1}{(m-x)^{2}}=\frac{(n-x)^{2}-(m-x)^{2}}{(n-x)^{2}(m-x)^{2}}=\frac{(n-m)((n-x)+(m-x))}{(n-x)^{2}(m-x)^{2}}\geq 0$. Thus $f$ is convex on $[0,m)$. This yields additional bounds: Jensen’s inequality gives
$f\left(\frac{t-1}{2}\right)=f\left(\frac{1}{t}\sum_{i=0}^{t-1}i\right)\leq\frac{1}{t}\sum_{i=0}^{t-1}f(i)\leq\frac{1}{t}\sum_{i=0}^{t-1}\frac{(t-1-i)f(0)+if(t-1)}{t-1}=\frac{f(0)+f(t-1)}{2},$ (B.14)
which rearranges to
$\left(\frac{n-\frac{t-1}{2}}{m-\frac{t-1}{2}}\right)^{t}\leq\frac{{n\choose t}}{{m\choose t}}=\exp\left(\sum_{i=0}^{t-1}f(i)\right)\leq\left(\frac{n}{m}\cdot\frac{n-t+1}{m-t+1}\right)^{t/2}.$ (B.15)
Several bounds in Table 6.1 are derived from Corollary 3.8. In particular, if $n-m=\frac{cn}{t+c}$, then
$C(n,m,t)$ $<\left(\frac{ne}{m}\right)^{t}\cdot\min\{1+m\log 2,1+t\log m\}+1$ (B.16)
$=e^{t}\left(1-\frac{n-m}{n}\right)^{-t}\cdot\min\{1+m\log 2,1+t\log m\}+1$
$=e^{t}\left(1-\frac{c}{t+c}\right)^{-t}\cdot\min\{1+m\log 2,1+t\log m\}+1$
$=e^{t}\left(1+\frac{c}{t}\right)^{t}\cdot\min\{1+m\log 2,1+t\log m\}+1$
$\leq e^{t+c}\cdot(1+t\log m)+1=O(e^{t+c}\cdot t\log m).$ ($1+x\leq e^{x}$)
Alternatively, Equation 3.15 gives
$C(n,m,t)=C(\ell(t+c),\ell t,t)\leq C(t+c,t,t)={t+c\choose t}\leq t^{c}.$ (B.17)
Setting $\ell=\frac{n}{t+c}\in\mathbb{Z}$ yields $n-m=\ell c=\frac{cn}{t+c}$.